OpenStudy (usukidoll):

Find the area of the region common to the interiors of the cardioids r = 1+cospheta and r = 1-cospheta

OpenStudy (usukidoll):

@zepdrix

zepdrix (zepdrix):

Oh you chose to split it into 4 regions? :) that's clever hehe

OpenStudy (usukidoll):

yeah no way in hexland that I would write a lot of integrals

OpenStudy (usukidoll):

it was sort of easy to see which came first...the one with the minus sign.

OpenStudy (usukidoll):

does it look ok???

zepdrix (zepdrix):

Mmm I'm having trouble getting through this one :( I would have approached it differently. Area of polar functions:\[\Large \frac{1}{2}\int\limits r^2\;d\theta\]

OpenStudy (usukidoll):

O_O

OpenStudy (anonymous):

That's what you get after one integration, so it shouldn't surprise anyone.

OpenStudy (anonymous):

Anyway, I'd say the lower limit of \(r(\theta)\) should be \(0\).

zepdrix (zepdrix):

^

OpenStudy (usukidoll):

oh now I'm seeing this... more like r = 0 for the lower one and the r upper should be 1 because you need to solve for the r's or something like 1+cospheta=1-cospheta.

zepdrix (zepdrix):

mm no the upper isn't 1. See how the upper boundary of the radius is changing? it's the function r=1-cos theta.

OpenStudy (anonymous):